# The generalized Hierarchical Gaussian Filter

### v2

@@ 3 Belief updates in the HGF: A network of nodes
   (re)compute $\mathrm{prediction}^{(k)}_{a}$
$\leftarrow$ receive $\mathrm{PE}^{(k)}_{child}$ from child node UPDATE step
@@
given: $\mathrm{PE}^{(k)}_{child}$ and $\mathrm{prediction}^{(k)}_{a}$rightarrow send $\mathrm{posterior}^{(k)}_{a}$ to child node PE step compute $\mathrm{PE}^{(k)}_{a}$ given: $\mathrm{prediction}^{(k)}_{a}$ and $\mathrm{posterior}^{(k)}_{a}$rightarrow send $\mathrm{PE}^{(k)}_{a}$ to parent node $\leftarrow$ receive $\mathrm{posterior}^{(k)}_{parent}$ from parent node PREDICTION step compute $\mathrm{prediction}^{(k+1)}_{a}$ given: $\mathrm{posterior}^{(k)}_{a}$ and $\mathrm{posterior}^{(k)}_{parent}$
@@
In the following, we examine the exact computations at time step $k$ within a node for each of the three steps introduced above. We will compare the relevant computations between volatility and value coupling and identify the messages that have to be sent and received in each step. Since the update step relies on quantities computed in the (previous) prediction step, we start with the computation of the predictions for the current time step (which, as explained above, we think of as being computed prior to the arrival of a new input). We will first only consider the case of linear value coupling (alongside volatility coupling), and then separately examine any differences in these computations for the case of nonlinear value coupling. The results of this analysis are summarized in Figures 3 – 5.
@@
The total predicted mean drift equals the sum of constant term $\rho_{a}$ (the tonic drift parameter of node $a$) and the sum of posterior means of all value parents $\mu_{b_{i}}^{(k-1)}$ (where parents are indexed by $i$) at $k-1$, weighted by their connection strengths $\alpha_{b_{i},a}$:
@@
where $t^{(k)}$ denotes the time that has passed between $k-1$ and $k$, and $N_{vapa}$ is the number of value parents. Similarly, the total predicted volatility $\Omega_{a}^{(k)}$ is a function of a constant term $\omega_{a}$ (the tonic volatility parameter) and the posterior means of all volatility parents $\mu_{\check{a}_{j}}^{(k-1)}$ at the previous time point $k-1$, weighted by their connection strengths $\kappa_{\check{a}_{j},a}$:
@@
This quantity summarizes the size of the belief update in node b due to changes in node a (before accounting for connection strength) and is often of interest in experimental investigations of neural correlates of prediction errors in belief updating (Iglesias et al., 2013a; Diaconescu et al., 2020; Weber et al., 2020, 2022).
@@
For a node $\check{a}$ which is the volatility parent of node $a$, the update equations for computing a new posterior mean $\mu_{\check{a}}^{(k)}$ and a new posterior precision $\pi_{\check{a}}^{(k)}$ have been described by Mathys et al. (2011). Here, we will introduce two changes to the notation to simplify the equations themselves and their implementation:
@@
$\begin{split}\Delta_{a}^{(k)}\equiv\delta_{a}^{(k,VOPE)}&:=\hat{\pi}_{a}^{(k)}\left(\frac{1}{\pi_{a}^{(k)}}+\left(\delta_{a}^{(k)}\right)^{2}\right)-1\\ &=\frac{\hat{\pi}_{a}^{(k)}}{\pi_{a}^{(k)}}+\hat{\pi}_{a}^{(k)}\left(\delta_{a}^{(k)}\right)^{2}-1.\\ \end{split}$ (42)
For a derivation of this definition based on the equations given in Mathys et al. (2011), cf.Appendix 6.2.
@@
An interesting difference to the implied message passing in predictive coding proposals (Bastos et al., 2012; Shipp, 2016) arises from the update step: The HGF architecture requires that not only (precision-weighted) prediction errors are being sent bottom-up between nodes, but also estimates of prediction precision ($\hat{\pi}_{a}$ or $\gamma_{a}$) which serve to update belief precision in the higher-level node.
@@
We have thus far focused our discussion on nodes that represent beliefs about continuous states which evolve in time as random walks (whether Gaussian or auto-regressive). However, the generalized HGF can also accommodate beliefs about any type of state governed by an exponential family distribution by filtering this distribution’s sufficient statistics Mathys & Weber (2020). This makes it possible to track binary and categorical states, where value parents track the probabilities with which one of several possible states is occupied, or with which transitions happen. This can be applied in a range of contexts, from experimental tasks where categorical probabilities must be learned Marshall et al. (2016) to inversions of other classes of discrete state space generative models like the Partially Observable Markov Decision Process models often used in Active Inference (Parr et al., 2022).
@@
We have presented the computations for these steps for the two different kinds of coupling that the HGF comprises: value coupling and volatility coupling. While the update equations for volatility coupling have been derived and discussed previously (Mathys et al., 2011, 2014a), approximately Bayes-optimal inference equations for (linear and nonlinear) value coupling under the HGF have not been considered prior to our treatment here. Furthermore, our analysis identifies not only the computations entailed by each computational step, but also the message passing between nodes that is required by each step. This is interesting from a theoretical point of view, where we can compare our architecture to other proposals of belief propagation.
From a practical point of view, the division of the belief updating machinery into subunits (nodes) allows for a modular implementation, where networks can easily be extended and modified by adding or removing nodes, or by changing the type of coupling between nodes, without having to derive the relevant equations for the whole network anew. In two open-source projects, we provide such an implementation (in Python: https://github.com/ilabcode/pyhgf, Legrand et al. (in prep); in Julia: https://github.com/ilabcode/HierarchicalGaussianFiltering.jl, Thestrup Waade et al. (in prep)), which allows users to flexibly design their own HGF structures that can be used for simulation and empirical parameter estimation. These tools will also be available as part of the TAPAS software collection (Frässle et al., 2021).
@@
4 Examples
In this section, we provide a few example simulations which show the range of generative models our generalization of the hgf encompasses. In the first example, two hidden states separately generate two streams of observations (Figure 6). Each state also has its own phasic volatility parent. However, these volatility states are both influenced by a shared value parent that encodes a higher-level state of global environmental volatility. The simulation shows that the inference network is able to infer the individual as well as the shared (“global”) volatility states. The level of noise in the outcomes ($u_{a}$ and $u_{b}$) was chosen relatively low. This model architecture could be used to explore how beliefs about environmental volatility generalize across different environmental states.
@@ 4 Examples
Second, we model a situation where an agent observes two cues (e.g., an auditory and a visual one) that each provide information about a hidden state of interest $x_{c}$ (such as the location of an object, Figure 7). Both observation sequences are subject to dynamically changing noise (hidden states $x_{\check{a}}$ and $x_{\check{b}}$). The simulations show that the agent can infer these changing noise levels and adjust its inference on the hidden state $x_{c}$ according to its current estimate of relative cue reliability.
Finally, in Figure 8, we consider an extension of a model that has found widespread application in decision neuroscience: Participants need to track the probability of a binary outcome over a sequence of observations (e.g., whether a certain choice or stimulus is rewarded or not). The sequence includes reversals and a stable period of $p=0.5$. In addition to the binary outcomes on every trial, the agent sometimes also has access to a (noisy) sample of the continuous hidden probability state. The simulation shows that inference on small jumps in probability as well as stable periods of $/50$ 50 is difficult based on binary observations alone, but that adding a continuous observation of the probability itself, even if it is only a very noisy readout of the actual probability, stabilizes inference.
@@
$\omega_{c}$ n/a 0 $\omega_{c}$3 -3 $\omega_{\check{a}}$ n/a -3
$\omega_{\check{a}}$3 -3 $\omega_{\check{a}}$ n/a -3 $\omega_{a}$ n/a -3
$\omega_{\check{b}}$3 -3 $\omega_{\check{b}}$ n/a -3 $\kappa_{a}$ 1 1
$\omega_{a}$2 -2 $\kappa_{a}$ 1 1 $\alpha_{b}$ 1 1
$\omega_{b}$2 -2 $\kappa_{b}$ 1 1 $\epsilon_{a}$ n/a -.5
$\alpha_{c,\check{a}}$ 2 .05 $\epsilon_{a}$3 -3 $x^{(0)}_{a}$ n/a (0, 1)
$\alpha_{c,\check{b}}$ 2 .05 $\epsilon_{b}$3 -3 $x^{(0)}_{\check{a}}$ n/a (4, 1)
$\kappa_{a}$5 .5 $x^{(0)}_{c}$ 0 (0, 1)
$\kappa_{b}$5 .5 $x^{(0)}_{\check{a}}$ n/a (0, 1)
@@ 5 Conclusions
Second, we present a modular architecture for HGF networks, where beliefs represent nodes that perform three basic computational steps: an Update step, a PE step, and a prediction step, in response to new input. While the equations for these steps differ depending on the coupling of a node to other nodes, we identify a generic structure that allows for a modular implementation, in which nodes can easily be added to or removed from a network, without having to derive the corresponding update equations for the model anew. This feature takes significant load off researchers wanting to apply this modelling framework and to create custom models that suit their experiments. We provide such implementations in two open-source projects (in Python: https://github.com/ilabcode/pyhgf, Legrand et al. (in prep); in Julia: https://github.com/ilabcode/HierarchicalGaussianFiltering.jl, Thestrup Waade et al. (in prep); inclusion in the TAPAS software collection (Frässle et al., 2021) is pending).
Finally, by considering the case of nonlinear value coupling and deriving the message passing scheme implied by this, we enable a formal comparison to other proposed architectures for hierarchical Bayesian inference, most prominently Bayesian (or generalized) predictive coding (Bastos et al., 2012; Shipp, 2016) and belief propagation in active inference (e.g., Fig. 5.1 in Parr et al., 2022).
@@ 5.1 Modelling different sources of uncertainty
Whenever agents are faced with observations that violate their expectations, they need to arbitrate between different explanations – has the world changed, requiring an update of beliefs about hidden states, or was the deviation merely due to noise in their observations? As has been shown previously (Mathys et al., 2011, 2014a), the HGF models belief updating in an agent who takes into account several forms of uncertainty for determining the optimal learning rate in the face of new observations: sensory uncertainty (how noisy are the sensory inputs I receive), informational uncertainty (how much do I already know about the hidden state that generates the inputs), and environmental uncertainty (what is the rate of change I expect in the hidden state). All of these together will determine whether (and how much) the agent updates its beliefs about a hidden state in response to unexpected observations.
@@
This flexibility in building models of hierarchically interacting states in the world allows for some particularly interesting use cases. In Figure 6, we have provided an example where two hidden states evolve with their own respective evolution rates (both determined by a tonic component and a phasic component), but share a higher-level value parent (“global volatility”) that drives changes in the mean of their respective phasic volatility states. This setup allows for separate estimation of ”local” volatility (specific to each hidden state) and more global influences on volatility. For example, the rate of change in the availability of two different foods in the environment might vary over time and seasons in a way that is specific to each type of food, but when switching to a different environment (or after a global change to the overall climate), the availability of both foods might become more or less stable. Investigating how the brain represents each form of volatility, and thus adjusts learning rates in a modality-specific as opposed to a general manner is an important part of understanding how the brain achieves and maintains the delicate balance of precision across different hierarchical levels (Kanai et al., 2015; Clark, 2013) which appears crucial for mental health (Petzschner et al., 2017; Sterzer et al., 2018).
Finally, agents are confronted not only with phasic variations of volatility (driving changes in the agent’s environmental uncertainty), but also with dynamically changing sensory (or observation) noise. While existing modelling approaches have focused on either accounting for changes in volatility (Behrens et al., 2007; Mathys et al., 2011; Piray & Daw, 2020) or changes in stochasticity or noise (Lee et al., 2020; Nassar et al., 2010), it has recently been pointed out that real-world agents need to able to detect (and distinguish) changes in both at the same time (Piray & Daw, 2021). In the HGF, dynamic changes in observation noise can be accommodated by hidden states that serve as noise parents to observable outcome states (see Appendix 6.3 for the equations, Figure 7 for an example simulation, and Mikus et al. (2023) for an application to a dataset). Jointly modelling an agent’s inference on volatility and stochasticity will be crucial to understanding the computational origin of maladaptive inference and learning in different psychiatric conditions (Piray & Daw, 2021; Pulcu & Browning, 2019; Mikus et al., 2023).

### v3

@@ 3 Belief updates in the HGF: A network of nodes
(re)compute $\mathrm{prediction}^{(k)}_{a} \leftarrow$ receive $\mathrm{PE}^{(k)}_{child}$ from child node UPDATE step
@@
given: $\mathrm{PE}^{(k)}_{child}$ and $\mathrm{prediction}^{(k)}_{a} \rightarrow$ send $\mathrm{posterior}^{(k)}_{a}$ to child node PE step compute $\mathrm{PE}^{(k)}_{a}$ given: $\mathrm{prediction}^{(k)}_{a}$ and $\mathrm{posterior}^{(k)}_{a} \rightarrow$ send $\mathrm{PE}^{(k)}_{a}$ to parent node $\leftarrow$ receive $\mathrm{posterior}^{(k)}_{parent}$ from parent node PREDICTION step compute $\mathrm{prediction}^{(k+1)}_{a}$ given: $\mathrm{posterior}^{(k)}_{a}$ and $\mathrm{posterior}^{(k)}_{parent}$
@@
In the following, we examine the exact computations at time step $k$ within a node for each of the three steps introduced above. We will compare the relevant computations between volatility and value coupling and identify the messages that have to be sent and received in each step. Since the update step relies on quantities computed in the (previous) prediction step, we start with the computation of the predictions for the current time step (which, as explained above, we think of as being computed prior to the arrival of a new input). We will first only consider the case of linear value coupling (alongside volatility coupling), and then separately examine any differences in these computations for the case of nonlinear value coupling. The results of this analysis are summarized in Figures 35.
@@
The total predicted mean drift equals the sum of constant term $\rho_{a}$ (the tonic drift parameter of node $a$) and the sum of posterior means of all value parents $\mu_{b_{i}}^{(k-1)}$ (where parents are indexed by $i$) at k1k-1, weighted by their connection strengths $\alpha_{b_{i},a}$:
@@
where $t^{(k)}$ denotes the time that has passed between k1k-1 and $k$, and $N_{vapa}$ is the number of value parents. Similarly, the total predicted volatility $\Omega_{a}^{(k)}$ is a function of a constant term $\omega_{a}$ (the tonic volatility parameter) and the posterior means of all volatility parents $\mu_{\check{a}_{j}}^{(k-1)}$ at the previous time point k1k-1, weighted by their connection strengths $\kappa_{\check{a}_{j},a}$:
@@
This quantity summarizes the size of the belief update in node b due to changes in node a (before accounting for connection strength) and is often of interest in experimental investigations of neural correlates of prediction errors in belief updating (iglesias_hierarchical_2013; diaconescu_hierarchical_2020; weber2020; weber2022).
@@
For a node $\check{a}$ which is the volatility parent of node $a$, the update equations for computing a new posterior mean $\mu_{\check{a}}^{(k)}$ and a new posterior precision $\pi_{\check{a}}^{(k)}$ have been described by mathys2011. Here, we will introduce two changes to the notation to simplify the equations themselves and their implementation:
@@
[…]
For a derivation of this definition based on the equations given in mathys2011, cf.Appendix 6.2.
@@
An interesting difference to the implied message passing in predictive coding proposals (bastos2012; shipp2016) arises from the update step: The HGF architecture requires that not only (precision-weighted) prediction errors are being sent bottom-up between nodes, but also estimates of prediction precision ($\hat{\pi}_{a}$ or $\gamma_{a}$) which serve to update belief precision in the higher-level node.
@@
We have thus far focused our discussion on nodes that represent beliefs about continuous states which evolve in time as random walks (whether Gaussian or auto-regressive). However, the generalized HGF can also accommodate beliefs about any type of state governed by an exponential family distribution by filtering this distribution’s sufficient statistics mathys_weber_2020. This makes it possible to track binary and categorical states, where value parents track the probabilities with which one of several possible states is occupied, or with which transitions happen. This can be applied in a range of contexts, from experimental tasks where categorical probabilities must be learned marshall_pharmacological_2016 to inversions of other classes of discrete state space generative models like the Partially Observable Markov Decision Process models often used in Active Inference (parr2022).
@@
We have presented the computations for these steps for the two different kinds of coupling that the HGF comprises: value coupling and volatility coupling. While the update equations for volatility coupling have been derived and discussed previously (mathys2011; mathys2014), approximately Bayes-optimal inference equations for (linear and nonlinear) value coupling under the HGF have not been considered prior to our treatment here. Furthermore, our analysis identifies not only the computations entailed by each computational step, but also the message passing between nodes that is required by each step. This is interesting from a theoretical point of view, where we can compare our architecture to other proposals of belief propagation.
From a practical point of view, the division of the belief updating machinery into subunits (nodes) allows for a modular implementation, where networks can easily be extended and modified by adding or removing nodes, or by changing the type of coupling between nodes, without having to derive the relevant equations for the whole network anew. In two open-source projects, we provide such an implementation (in Python and Julia, legrand2024; waade_in_prep_2023), which allows users to flexibly design their own HGF structures that can be used for simulation and empirical parameter estimation. These tools are freely available as part of the TAPAS software collection (frassle_tapas_2021, https://github.com/ComputationalPsychiatry).
@@
4 Practical applications
In this section, we provide a few example simulations which show the range of generative models our generalization of the hgf encompasses. We focus on relatively low-dimensional situations that have been of interest in empirical research in cognitive neuroscience and computational psychiatry, but have so far been challenging to model.
@@ 4 Practical applications
In the first example, we exploit the fact that nodes in the generalised HGF can share parent nodes. This allows us to model generalisation. For example, an agent who experiences high volatility in one context or domain might change their higher-level beliefs about the stability of the world more generally, and thus start to treat the dynamics in other domains or contexts as similarly volatile. In our minimal example for this, two hidden states separately generate two streams of observations (Figure 6). These could for example represent the observed behaviours of two different acquaintances the agent interacts with, or the availability of food in two different patches. Each state also has its own phasic volatility parent. However, these volatility states are both influenced by a shared value parent that encodes a higher-level state of global environmental volatility. For example, while the availability of food varies independently over time in the two patches, the overall season might produce upticks in volatility across different food sources simultaneously. We simulate two hidden states ($x_{a}$ and $x_{b}$), with one of them more volatile than the other (volatility parent $x_{\check{a}}>x_{\check{b}}$). The shared value (drift) parent of the two volatility nodes, $x_{c}$, produces a marked up-tick (fast drift starting at trial 50) and one down-tick in the individual volatilities. The simulation shows that under the chosen parameter settings the inference network is able to infer the individual as well as the shared (“global”) volatility states. The level of noise in the outcomes ($u_{a}$ and $u_{b}$) was chosen relatively low. This model architecture could be used to explore how beliefs about environmental volatility generalize across different environmental states, and under which conditions this breaks down (e.g., overgeneralisation).
@@
In the second example, we revisit a widely studied perceptual process: multi-sensory cue combination. When humans observe two cues (e.g., an auditory and a visual one) that each provide information about a hidden state of interest $x_{c}$ (such as the location of an object, Figure 7), they can integrate both sources of information, weighing each source according to its reliability (ernst2002). Under natural circumstances, these reliabilities might dynamically change, for example due to physiological fluctuations or changes in the external environment. Here, we model a situation where two observation sequences are each subject to dynamically changing noise (hidden states $x_{\check{a}}$ and $x_{\check{b}}$) using nodes connected via noise coupling. Our model features a shared value parent node (hidden state $x_{c}$) that generates both outcomes. The simulations show that the agent can infer these changing noise levels and adjust its inference on the shared hidden state $x_{c}$ according to its current estimate of relative cue reliability. Critically, this setup allows us to model and infer changes in the subjective beliefs about the relative reliabilities and their impact on sequential belief updating. The states $x_{\check{a}}$ and $x_{\check{b}}$ could additionally be children of higher-level states, for example, to model higher-level beliefs about one’s confidence in different sensory modalities. The model architecture is relevant for cognitive tasks that assess perceptual integration between different modalities (such as integrating interoceptive signals with auditory feedback when inferring one’s own heart rate) in trial-by-trial fashion. Maladaptive over- or under-weighting of interoceptive versus exteroceptive signals is believed to drive symptoms of mental health (paulus2010).
Finally, in our last example, we exploit the fact that the same node in the generalised HGF can have parent or child nodes with different coupling types, and that any node within a network can be connected with an outcome node, if the state becomes somehow observable. In Figure 8, we consider an extension of an experiment that has found widespread application in decision neuroscience: a reversal learning task, where participants track the probability of a binary outcome over a sequence of observations (e.g., whether a certain choice or stimulus is rewarded or not). The example sequence chosen here includes probability reversals and stable periods of $p=0.5$ (i.e., both outcomes are equally likely). The typical model for this is that the participant infers on a hidden state $x_{a}$ (and potentially its volatility $x_{\check{a}}$) which represents the probability or tendency for one outcome over the other. In our example, we add an additional (continuous) observation to this hidden state $x_{a}$, in other words, in addition to the binary outcomes on every trial, the agent sometimes also has access to a (noisy) sample of the continuous hidden probability state. The simulation shows that inference on small jumps in probability as well as stable periods of $50/50$ is difficult based on binary observations alone (trials 50 to 150), but that adding a continuous observation of the probability itself, even if it is only a very noisy readout of the actual probability, stabilizes inference. The model architecture thus allows us to make predictions about the relative usefulness of different types of observations when inferring on a hidden state.
@@
$\omega_{c}$ n/a 0 $\omega_{c}$ -3 -3 $\omega_{\check{a}}$ n/a -3
$\omega_{\check{a}}$ -3 -3 $\omega_{\check{a}}$ n/a -3 $\omega_{a}$ n/a -3
$\omega_{\check{b}}$ -3 -3 $\omega_{\check{b}}$ n/a -3 $\kappa_{a}$ 1 1
$\omega_{a}$ -2 -2 $\kappa_{a}$ 1 1 $\alpha_{b}$ 1 1
$\omega_{b}$ -2 -2 $\kappa_{b}$ 1 1 $\epsilon_{a}$ n/a -.5
$\alpha_{c,\check{a}}$ 2 .05 $\epsilon_{a}$ -3 -3 $x^{(0)}_{a}$ n/a (0, 1)
$\alpha_{c,\check{b}}$ 2 .05 $\epsilon_{b}$ -3 -3 $x^{(0)}_{\check{a}}$ n/a (4, 1)
$\kappa_{a}$ .5 .5 $x^{(0)}_{c}$ 0 (0, 1)
$\kappa_{b}$ .5 .5 $x^{(0)}_{\check{a}}$ n/a (0, 1)
@@ 5 Conclusions
Second, we present a modular architecture for HGF networks, where beliefs represent nodes that perform three basic computational steps: an Update step, a PE step, and a prediction step, in response to new input. While the equations for these steps differ depending on the coupling of a node to other nodes, we identify a generic structure that allows for a modular implementation, in which nodes can easily be added to or removed from a network, without having to derive the corresponding update equations for the model anew. This feature takes significant load off researchers wanting to apply this modelling framework and to create custom models that suit their experiments. We provide such implementations in two open-source projects (legrand2024; waade_in_prep_2023) which are available as part of the TAPAS software collection (https://github.com/ComputationalPsychiatry frassle_tapas_2021).
Finally, by considering the case of nonlinear value coupling and deriving the message passing scheme implied by this, we enable a formal comparison to other proposed architectures for hierarchical Bayesian inference, most prominently Bayesian (or generalized) predictive coding (bastos2012; shipp2016) and belief propagation in active inference (e.g., Fig. 5.1 in parr2022).
@@ 5.1 Modelling different sources of uncertainty
Whenever agents are faced with observations that violate their expectations, they need to arbitrate between different explanations – has the world changed, requiring an update of beliefs about hidden states, or was the deviation merely due to noise in their observations? As has been shown previously (mathys2011; mathys2014), the HGF models belief updating in an agent who takes into account several forms of uncertainty for determining the optimal learning rate in the face of new observations: sensory uncertainty (how noisy are the sensory inputs I receive), informational uncertainty (how much do I already know about the hidden state that generates the inputs), and environmental uncertainty (what is the rate of change I expect in the hidden state). All of these together will determine whether (and how much) the agent updates its beliefs about a hidden state in response to unexpected observations.
@@
This flexibility in building models of hierarchically interacting states in the world allows for some particularly interesting use cases. In Figure 6, we have provided an example where two hidden states evolve with their own respective evolution rates (both determined by a tonic component and a phasic component), but share a higher-level value parent (“global volatility”) that drives changes in the mean of their respective phasic volatility states. This setup allows for separate estimation of ”local” volatility (specific to each hidden state) and more global influences on volatility. For example, the rate of change in the availability of two different foods in the environment might vary over time and seasons in a way that is specific to each type of food, but when switching to a different environment (or after a global change to the overall climate), the availability of both foods might become more or less stable. Investigating how the brain represents each form of volatility, and thus adjusts learning rates in a modality-specific as opposed to a general manner is an important part of understanding how the brain achieves and maintains the delicate balance of precision across different hierarchical levels (kanai2015; clark2013) which appears crucial for mental health (petzschner2017; sterzer2018).
Finally, agents are confronted not only with phasic variations of volatility (driving changes in the agent’s environmental uncertainty), but also with dynamically changing sensory (or observation) noise. While existing modelling approaches have focused on either accounting for changes in volatility (behrens2007; mathys2011; piray2020a) or block-wise changes in stochasticity or noise (lee2020; nassar2010; nassar2012; nassar2016), it has recently been pointed out that real-world agents need to able to detect (and distinguish) dynamic changes in both at the same time (piray2020b), and recent empirical work has demonstrated that human participants are indeed able to do so (piray2024; foucault2025). In the HGF, dynamic changes in observation noise can easily be accommodated by hidden states that serve as noise parents to observable outcome states (see Appendix 6.3 for the equations, Figure 7 for an example simulation, and mikus2025, for an application to a dataset in computational psychiatry).